### v2

# Comparing Pre-trained Human Language Models: Is it Better with Human Context as Groups, Individual Traits, or Both?

Abstract
Incorporating human context into language models is the next frontier for human-centered natural language processing. Currently, two pre-training methods exist: group-wise attributes (e.g., over-45-year-olds) or individual traits. Group attributes are coarse — not all 45-year-olds write the same way — while modeling individual traits allows for a more personalized representation, but requires more complex modeling and data. So far, it is unclear which pre-training approach benefits what tasks. We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on 5 user- and document-level tasks. We find that pre-training with both group and individual features significantly improves the two user-level regression tasks like age estimation and personality assessment. Pre-training on individual users significantly improves the three document-level classification tasks like stance and topic detection. It even does well for downstream tasks without historical user data. Our results suggest both approaches have specific use cases, opening new avenues for human-centered language modeling.
Comparing Pre-trained Human Language Models:
Is it Better with Human Context as Groups, Individual Traits, or Both?
Nikita Soni1, Niranjan Balasubramanian1, H. Andrew Schwartz1,  and Dirk Hovy2 1Stony Brook University, 2Bocconi University {nisoni, niranjan, has}@cs.stonybrook.edu, dirk.hovy@unibocconi.it
1 Introduction
Language varies between people. To capture this notion, two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language. The first approach focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group. These socio-demographic attributes include age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more. The second approach focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
Figure 1: Pre-training a language model with no human context, with socio-demographic group attribute, with individual traits, and with both group and individual traits.
While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2023): People are not defined by their group membership alone Orlikowski et al. (2023), but individual traits might not generalize. We compare the downstream performance of models pre-trained with the two approaches as well as their combination. We answer the following broader research questions:
(RQ1): How can we incorporate group and/or individual human context into pre-training?
(RQ2): How does pre-training with different human contexts differ in terms of downstream performance for different tasks?
We use the monolingual socio-demographically adapted model from Hung et al. (2023) and the HaRT model from Soni et al. (2022) for group and individual human contexts, respectively. For a model trained with both individual and group human context, we use a multi-task learning PLM based on HaRT to create a model, GRIT, in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness). We test all systems on five user- and document-level tasks.
Note that because we focus on empirically comparing pre-training group and individual traits in PLMs, we cannot compare to large language models like GPT4, which do not support stratification to either attribute via pre-training. Recent studies have explored methodologies like user-adapters (Zhong et al., 2021), and user-centric prompting (Li et al., 2023) to include human context into models. In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts.
We find that PLM pre-trained on individuals and group features enhances 2 user-level regression tasks: age estimation and personality assessment from multiple user documents. These user-level tasks focus on individual people. Our findings, that they are best modeled as a combination of their group attributes and unique characteristics, conforms with the notion that a person is a mix of both. Document-level categorization tasks, like stance detection, are more personal than group based. A PLM pre-trained on individual human context alone improves 3 document-level classification tasks.
By their very nature, models of this kind touch upon sensitive user information. For this reason, we take a responsible release strategy, making only the code for the comparisons publicly available and the exact splits of the TrustPilot and Stance datasets used. We build on top of the publicly available code for HaRT and Hung et al. (2023). We acquired the model and data in a secure manner from the authors of Soni et al. (2022) and TrustPilot data splits from the authors of Hung et al. (2023). For more information about the model and data, see Sections 3 and 4. For a discussion of the ethical implications of the model and data, see Ethical Considerations section.
Contributions.
(1) We compare pre-training several language models with individual users, group socio-demographic features, and both traits.
(2) We evaluate the three specific pre-training strategies on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category prediction).
(3) We find that the two user-level regression tasks perform better when pre-trained with both strategies (GRIT) and the three single document-level tasks perform better on fine-tuning a language model pre-trained with individual context alone (HaRT).
2 Integrating Human Context in PLMs
For our comparison, we use three systems representing the paradigms of pre-training with human context (Figure 1). We want to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts. As noted earlier, we focus on recent approaches for pre-training language models with the additional human context.
Pre-training with group context.
We build on Hung et al. (2023), a recent model to explore demographic adaptation in transformer-based PLMs. They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author. They use the multilingual reviews dataset with demographic labels from Trustpilot [^1] https://www.trustpilot.com/ Hovy (2015). They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection. We use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4). We compare to the results from Hung et al. (2023) with monolingual BERT on a US-English dataset with out-of-domain demographic (age) specialization. Out-of-domain data is the Blogs authorship corpus Schler et al. (2006), and in-domain means Trustpilot corpus. To be consistent and fair in comparing with other human context pre-training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
Pre-training with individual human context.
Soni et al. (2022) introduced human language modeling (HuLM) in pre-trained LMs. I.e., pre-training a regular language modeling but including a dynamic individual human context vector derived from the authors’ historical texts. This vector captures the human states in which the text was generated. It also adds coherence to texts generated by the same author. They introduce a Human-aware Recurrent Transformer (HaRT), an autoregressive PLM for human language modeling. They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks. We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
Pre-training with both group and individual human context.
We train a PLM to integrate both individual and group human context in language modeling pre-training. We extend HaRT by predicting an additional human group attribute in a dynamic muti-task learning setup similar to Hung et al.. To facilitate comparison, we combine the pre-training inducing the individual human context through the author’s language and inject the group context by predicting a group attribute of the author.
We choose downstream tasks from both prior works (Soni et al., 2022; Hung et al., 2023) to compare our three language models pre-trained with different levels of human contexts: two user-level regression tasks and three single document-level classification task.
3 Models
3.1 Pre-training with individual human context
HaRT.
Soni et al. (2022) introduced HaRT to incorporate individual human context into PLMs. They use a 12-layered autoregressive GPT-2 based architecture with a modified self-attention computation at layer 2. This modification to the query vector now includes the individual human context via a dynamic user-state vector.
[…]
where $IN$ is the insert layer (layer 2), $Q_{i}$ is the query vector under computation, $H_{i}$ is the hidden states vector, and $U_{i-1}$ is the user-state vector derived from the previous block of language seen from the user. All the text from a user is processed in the same forward pass with recurrent processing of blocks of fixed-length (1024) tokens chunked after temporally ordering the social media posts by created time. The user state is recurrently updated using the hidden states from layer 11 and computed as follows:
$U_{i}=tanh(W_{U}U_{i-1}+W_{H}H^{(E)})$
where, $E$ is the extract layer (layer 11), $U_{i}$ is the updated user-state vector, $U_{i-1}$ is the user-state vector from the previous block, and $H^{E}$ is the hidden states vector from layer 11. This formulation of updating the user-state vector extends the previous user-state vector information with the current language block’s information.
HuLM pre-training task.
HaRT is pre-trained for the human language modeling (HuLM) task defined as predicting the next token given the previous tokens while conditioning on previous user state $U_{1:t-1}$ (Soni et al., 2022) .
$Pr(\mathbf{W}_{t}|\mathbf{U}_{t-1})=\prod_{i=1}^{n}Pr(w_{t,i}|w_{t,1:i-1},\mathbf{U}_{1:t-1})$
This is translated into a pre-training objective to maximize:
$\prod_{a\in\textup{Users}}\prod_{t=1}^{|\mathcal{B}_{a}|}\prod_{i=1}^{|B_{t}^{(a)}|}Pr(w_{t,i}|w_{t,1:i-1},B_{1:t-1}^{(a)})$
where, $w_{t,i}$ is the $i^{th}$ token in the $t^{th}$ block ($B_{t}^{(a)}$) for user $a$. The tokens from the previous blocks are represented using HaRT’s recurrently updated user-state vector. Soni et al. use cross-entropy loss for the HuLM objective.
3.2 Pre-training with group human context
BERTDS and BERTage-MLM.
Hung et al. (2023) explore socio-demographic adapted BERT models to inject group human context into PLMs. We use the names BERTDS and BERTage-MLM to denote their demographic (age) specialization using the multi-task learning setup and demographic adaptation with masked language modeling respectively.
Multi-Task Learning.
Hung et al. (2023) train for domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, they adopt a dynamic MTL objective for training with group human context. Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions. This approach accounts for the different scales across multiple losses.
$\tilde{L_{t}}=\frac{1}{2\sigma_{mse}^{2}}L_{t}+\log\sigma_{t}$
Hung et al. minimize the sum of both the uncertainty adjusted losses: $\tilde{L_{mlm}}+\tilde{L_{dem}}$.
3.3 Pre-training with both individual and group human context
GRIT.
GRoup and Individual HaRT is a LM pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup. The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
$Pr(attribute|\overline{\mathbf{U}})$
Because of compute limitations, we chunk a user’s language history into blocks and process them in a single forward pass. Each block of text from a user results in a user-state vector. We use the average of the user-state vectors from each non-padded block of texts from an author to compute their final user-state representation. This representation is layer-normed and linearly transformed before making a continuous-valued prediction for the specific attribute.
We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope), respectively. The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$). Similar to Hung et al. (2023), we account for the homoscedastic uncertainty Kendall et al. (2018) of losses. We use the joint loss for a continuous and discrete output as derived in Kendall et al. (2018) and compute our joint objective as follows:
$\frac{1}{\sigma_{ce}^{2}}L_{ce}+\frac{1}{2\sigma_{mse}^{2}}L_{mse}+\log\sigma_{ce}+\log\sigma_{mse}$
where, $\sigma_{ce}$ and $\sigma_{mse}$ are the variances of the task-specific losses.
In practice, we use log variance for numerical stability and use the adjusted loss calculation as follows:
$\exp^{-\eta_{ce}}L_{ce}+\eta_{ce}+\frac{1}{2}(\exp^{-\eta_{mse}}L_{mse}+\eta_{mse})$
where $\eta_{x}=\log\sigma_{x}^{2}$. We let $\sigma_{ce}$ and $\sigma_{mse}$ be learnable parameters for the model. We do not halve the log-term of the cross-entropy loss since we found it to perform better with our multi-task learning experiments.
Pre-training data.
We use a subset of the pre-training data for HaRT, consisting of the demographics and personality information. This subset contains the Facebook posts from Park et al. (2015) as used by Soni et al.. Our dataset is consistent with the inclusion criteria for HaRT to ensure moderate language history for each user: we include English posts from users with at least 50 total posts and at least 1000 words. This dataset consists of just over 63,000 unique users, which we split into a training dataset consisting of messages from 56,930 users, a development dataset that consists of messages from 1836 users that were not part of the training set, and a test set of messages from a separate set of 4438 users that are neither in training nor the development set. To evaluate the human attribute prediction in GRITope, we use a subset of the test set consisting of messages from 1745 users to accommodate for questionnaire reliability. We use the Facebook posts for the HuLM task and the demographic and personality scores of consenting Facebook users (Kosinski et al., 2013) for the human attribute prediction task.
Training.
We use HaRT’s pre-trained weights as the base weights for GRIT and randomly initialize the newly introduced weights for human attribute prediction. GRIT is trained on our pre-training dataset using the 5e-5 learning rate after experimenting with a few learning rates, including that used for HaRT’s pre-training. Following HaRT, and due to computing limitations, each training instance is capped to 8 blocks of 1024 tokens each, with train batch size as 1 per device and evaluation batch size as 20 per device, trained over 2 GPUs for eight epochs. We explored multiple joint losses before resorting to the homoscedastic loss computation. Since HaRT caps to 4 train blocks for user-level downstream tasks, we also pre-train GRITage and GRITope with four training blocks.
3.4 Fine-Tuning
We fine-tune GRIT and HaRT for downstream document-level tasks. Each downstream task has a separate fine-tuned model that is initialized with the respective model’s pre-trained parameters and trained using the respective downstream task labels and an appropriate loss function. We also use the historical language of a user where available for any of the downstream tasks. We use the last predicted token’s representation to predict the label in document-level classification tasks. We experimented with fine-tuning GRIT for user-level regression tasks in multiple ways, including 1) similar to HaRT, by using the averaged user-state vectors from GRIT, 2) same as previous but fine-tuning only the history module, attribute prediction module, and the downstream task head, 3) freezing all the parameters of GRIT and fine-tuning the human attribute prediction module alone. However, we found continued training as described in section 3.5 to perform best.
We used the Optuna framework Akiba et al. (2019) for hyperparameter search, closely following the experimental settings in Soni et al..
3.5 Transfer Learning
We experiment with continuing pre-training GRIT models for each group attribute. To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness). We do the same for GRITope and continued MTL with predicting age.
4 Experiments
Our research goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined. To this end, we evaluate performances of the models defined in Section 3 on two user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification. We also compare against GPT-2HLC from Soni et al. (2022) as a PLM adapted to the social media domain but devoid of human context.
4.1 User Level Regression Tasks
We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment Kosinski et al. (2013); Park et al. (2015). These social scientific tasks require predicting continuous outcomes for a user given multiple documents written by them. We use the same data splits as used by Soni et al. (2022) for evaluating HaRT and GPT-2HLC
We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively. We further evaluate these models on the test sets for personality assessment and age estimation, respectively, after continuing training for these tasks, as described in section 3.5.
We use the results from Soni et al. (2022) for HaRT and GPT-2HLC which are directly comparable to GRIT models trained on the same data splits and metrics. Soni et al. fine-tuned the recurrence module of the pre-trained HaRT model for the tasks of age estimation and personality assessment using the average of user-states from non-padded blocks of texts from an author, resulting in two fine-tuned models. Similarly, they fine-tune the last two layers of the pre-trained GPT-2HLC model for these tasks. Since GPT-2HLC can not handle all text from a user in one pass, they average the predictions across all user messages corresponding to the same label for each message.
4.2 Document-Level Classification Tasks
We compare different models for stance detection vs. topic detection and age classification. These tasks classify a single input document (tweet message or a review) a user writes into label categories. For stance detection, we also use the historical messages of a user where available, as in Soni et al. (2022). We do not have the user information or any user historical language available for the other two tasks, so we evaluate on the single document input.
All models process the input document(s) and feed the layer-normed last non-padded token representation to the classification layer to classify the document into label categories. Only GRIT and HaRT incorporate user information and the historical language (where available). The other two models can only use the input document without a hierarchical structure to make the predictions. We compare with the results from Soni et al. (2022) and Hung et al. (2023) wherever applicable and fine-tune all the parameters of the respective pre-trained models and the classification heads for other task-model combinations using the standard cross-entropy loss.
Stance Detection
Given a single annotated tweet, this task predicts a user’s stance as in favor of, against, or neutral towards one of the five targets: atheism, climate change as a real concern, feminism, Hillary Clinton, and legalization of abortion. We fine-tune the models under comparison for each target separately. We report average of weighted F1 scores with three labels across all five targets. We use Soni et al. (2022)’s train/dev/test split over SemEval 2016 dataset (Mohammad et al., 2016). HaRT and GRIT models maintain the temporal accuracy by using only the messages posted earlier than the labeled messages from the extended dataset (Lynn et al., 2019) as a user’s historical language. We compare the results of fine-tuned GPT-2HLC, HaRT (Soni et al., 2022), and fine-tuned GRITage and GRITope.
Topic Detection
We use the US subset of the TrustPilot reviews dataset Hovy (2015) from two age groups: below 35 or above 45  [^2] As suggested by Hovy (2015), this split of the age ranges results in roughly equally-sized data sets and is non-contiguous, avoiding fuzzy boundaries.. Given a single review, this task predicts the review topics from five categories: Flights, Online marketplace, Fitness & Nutrition, Electronics, and Hotels. We use the same train, development, and test set splits as Hung et al. (2023) to eliminate any skew in the demographically-conditioned label distribution. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
Demographic Attribute Classification
We use the same subset of the TrustPilot dataset as for topic detection and the same train, development, and test splits from Hung et al. (2023). Given a single review, this task predicts the age group binary label (<35 years old or >45 years old). Age categories are equally represented in each set. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
4.3 Human Language Modeling
To compare the effects of individual and group factors on language modeling performance, we evaluate on the test set from the pre-trained data splits. We report and compare perplexity scores from the pre-trained GPT-2 (GPT-2frozen), GPT-2HLC, HaRT, GRITage and GRITope for the human language modeling task.
5 Results and Discussion
We report results for all the tasks here, discussing their respective impacts from pre-training LMs with individual human context, group context, and both individual and group context.
5.1 Comparisons Study
User-Level Regression Tasks.
Table 1 shows the two user-level regression task results. We find that just the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality. Additionally, comparing the transfer learning (Section 3.5) results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find the GRIT models to fare better.
Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it is still devoid of human context. HaRT adds pre-training with the individual human context, and GRIT adds pre-training with both group and individual human contexts (Figure 1). As Table 1 shows, there are gains going from GPT-2HLC (no human context) to HaRT (individual human context) and further to GRIT (individual + group human context). This indicates that pre-training PLMs with individual and group human context may benefit multi-document user-level regression tasks like the ones we considered.
Model Human Context Age ($r$) OPE ($r_{dis}$)
GPT-2HLC None 0.839 0.521
HaRT Individual 0.868 0.619
GRITage Ind + Grp 0.890 0.658
GRITope Ind + Grp $0.884$ $0.643$
Table 1: Pearson $r$ for age, disattenuated Pearson $r$ for openness. Pre-training with individual plus group context show benefits in estimating age and assessing personality (openness). Bold = best in column. We find no statistical difference between GRITage and GRITope for the task of age estimation. All other results show statistical significance $p<0.05$ using paired t-test.
Document-Level Classification Tasks.
Model Human Context Stance (F1wtd) TD (<35) (F1mac) TD (>45) (F1mac) AC (F1mac)
GPT-2HLC None 68.6 69.8 65.4 63.9
BERTage-MLM Group - 68.4 64.6 $61.9$
BERTDS Group - 69.3 65.0 $64.1$
HaRT Individual $\textbf{71.1}^{*}$ $\textbf{69.8}^{*}$ 65.6 $\textbf{64.3}^{*}$
GRITage Ind + Grp 70.8 69.2 64.5 62.7
GRITope Ind + Grp 70.1 66.5 64.8 61.2
[…]
Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute classification (AC). We see that task fine-tuned HaRT (individual human context) models performs better on all tasks..
HaRT models inherently have an additional context of the individual user and do not treat all inputs as if written by the same user. The tasks considered here relate more to personal opinions and preferences, rather than group-level ones. HaRT model is well-suited for incorporating such personalization due to its pre-training with individual human context. Additionally, the results indicate that the models pre-trained with group context (BERTDS) fare well in the group-based tasks of topic detection and age classification. Whereas, the models pre-trained on both individual and group human context (GRIT) appears to bring in noise for the group-based and personal stance detection tasks resulting in a slightly worse performance.
Further, it is important to note that the individual human context (HaRT) derived from the historical tweets from the users in the stance detection dataset provides for a richer human context. Although, the performance is not greatly hurt even if historical language is not available for certain tasks (TD and AC).
Perplexity.
We also compare the language modeling capability of the various models. Table 3 reports perplexity on the test set of 4438 users. The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT (individual human context) models perform best while GRIT (individual + group human context) models result in a slightly lower perplexity than HaRT. GRIT models are pre-trained under a multi-task learning setup which most likely is hurting the individual task performances, thus resulting in a slight dip in the perplexity compared to HaRT. Further, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al. (2022).
Model Human Context Test ($ppl$)
GPT-2frozen None 114.82
GPT-2HLC None 36.39
HaRT Individual 28.24
GRITage Ind + Grp 31.77
GRITope Ind + Grp 30.32
Table 3: Comparing perplexity on language modeling for models trained with individual and group contexts.
5.2 Error Analysis and Disparity
We conduct an error analysis as a function of a socio-demographic group attribute. Specifically, we measure GRIT and HaRT for error disparity (Shah et al., 2020) – – a systematic difference in error as a function of demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015). We analyzed outcome and error disparity on age and openness prediction from both the models with individual context (HaRT) and combined individual plus group context (GRIT).
First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set. Then, we compare the performance of our models across these buckets. Results from table 4 indicate that pre-training with individual and group context together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age. We see similar trends for assessing openness personality (Appendix Tables 6, 8). This suggests that the group attribute prediction maybe acting as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, that is averaged by the number of pairs (Shah et al., 2020). A lower averaged sum of differences implies lesser errors as a function of the age groups. Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 1113 0.223 0.394 0.393
18-21 1387 0.230 0.278 0.276
21-30 1557 0.512 0.531 0.519
30-45 695 0.485 0.530 0.520
45+ 248 0.106 0.205 0.180
Table 4: Pearson $r$ for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row. We find no statistical difference between GRITage and GRITope for buckets 21-30 and 30-45. All other results show statistical significance $p<0.05$ using paired t-test.
[…]
Age ($r$) 0.215 0.181 0.185
OPE ($r_{dis}$) 0.075 0.090 0.072
Table 5: Mean error disparity for age estimation and openness personality assessment over five age buckets. Bold indicates best in column (lower is better).
6 Related Work
Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models. Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups. These works improved sentiment analysis, polarity classification, and topic detection. Other researchers explored adapting to user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) to improve downstream tasks like sarcasm detection, and toxic language detection. Some studies go beyond explicit groups and learning individual representations latently Jaech and Ostendorf (2018); Delasalles et al. (2019) or via historical language Matero et al. (2021).
With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks. More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context. However, few studies have explored including human context within the pre-training regime of LMs. Hung et al. (2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup. Further, Soni et al. (2022) pre-train a LM with individual human context derived from user language. Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
7 Conclusion
NLP benefits from modeling latent human context, such as sociodemographic group traits or individual tendencies. A recent development has been to incorporate this additional human context into the PLMs’ pre-training regimen. However, humans exhibit varying degrees of group and individual characteristics. Learning about the impacts of pre-training with different types of human context will propel us forward in including human context in our base LLMs (Soni et al., 2023). To assess its impacts, we compare three types of PLMs pre-trained with sociodemographic group attributes, individual human contexts, or a combination across five user- and document-level tasks. We find that pre-training with individual and group human context improves the two user-level tasks: age and personality prediction. Pre-training with individual human context improves the three single-document classification tasks, including stance and topic detection. Pre-training solely on group context helps group-based document classification tasks, though not optimally. Our findings show a promising step towards modeling human context and provide valuable insights for the NLP community to think of additional strategies for improving models with task-dependent human context in pre-training.
Limitations
The purpose of our study is to compare the impacts of modeling sociodemographic group attributes and modeling individual user traits, and we use relevant models to represent each of the approaches. There are likely to be other ways to model these approaches and the models we use are only one of the ways. Additionally, these models in themselves have limitations like the blocks mechanism to process all the text from author induces compute requirements resulting in a capping of the number of blocks used for training. While it is also unclear how many blocks are sufficient to capture the human context, and if it is helpful to use the earliest language or the most recently used language in the capped number of blocks.
Secondly, some of the datasets (TrustPilot) used do not have appropriate user identification or historical language to create an individual human context.
Lastly, as noted earlier, models and data that touch upon sensitive user information require an extremely responsible usage and limit researchers to make them publicly available.
Ethical Considerations
Models that incorporate sociodemographic information need to be considered with special scrutiny. On the one hand, they have the potential to produce fairer and more inclusive results, because they can account for human language variation. On the other hand, they risk revealing identifying or sensitive information, which can lead to profiling and stereotyping. These may present opportunities for unintended malicious exploitations. For example, models that improve demographic groups prediction or psychological assessments could be used for targeting content for individuals without their awareness or consent. Such models may also risk release of private information of the research participant if trained on private data unchecked for exposing identifying information. For this reason, we take a conservative release strategy. While we support open research and reproducibility, data and privacy protection take precedence. Thus, we will only be releasing the code for our comparison study and the data that does not contain sensitive information i.e., stance detection datasets and TrustPilot datasets for topic detection and attribute classification. This is also in accordance with the DUA we have received from the authors of the papers/models that we employ in our work.
Our comparison study aims to guide and further speed the growing body of human-centered AI research. The models under comparison aim to enable applicability in the interdisciplinary studies of the human condition leading to helpful tools for psychological health. However, at this point these models are not intended for use in practice and should be evaluated for failures. All user-level tasks presented here were reviewed and approved or exempted by an academic institutional review board (IRB). Our studies are limited to US-English due to comparability reasons. However, similar effects are likely to hold for other languages, and should be evaluated in future work.
Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #585968. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Appendix
A.1 Experimental Settings
We closely follow the experimental settings from Soni et al. (2022) and similarly use Optuna framework (Akiba et al., 2019) for hyperparameter search. We search for learning rates between 5e-6 and 5e-4, and between 1e-7 and 1e-5 for different tasks. We will make our best found hyperparameter values publicly available with our code and results in the github repository. All experiments are run on NVIDIA RTX A6000 GPUs of 48GB. Pre-training takes approx 14 hours for 1 epoch and fine-tuning takes approx 1-4 hours depending on the task.
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 503 0.627 0.644 0.618
18-21 560 0.557 0.608 0.592
21-30 563 0.715 0.741 0.738
30-45 249 0.594 0.669 0.667
45+ 68 0.567 0.546 0.599
Table 6: Disattenuated pearson $r$ for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row.
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 1113 4.07 2.52 2.82
18-21 1387 6.52 4.00 3.89
21-30 1557 17.82 12.64 13.11
30-45 695 48.59 39.79 40.43
45+ 248 114.92 121.66 134.72
Table 7: Mean squared error for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 503 0.423 0.410 0.429
18-21 560 0.496 0.487 0.506
21-30 563 0.429 0.380 0.381
30-45 249 0.578 0.489 0.489
45+ 68 0.584 0.501 0.467
Table 8: Mean squared error for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).